OpenStudy (anonymous):

a parallelogram is a quadrilateral with two pairs of parallel sides , show that ABCD is a parallelogram if the coordinates of the vertices are A(-2,3) , B(2,7), C(8,5),and D(4,1)

OpenStudy (anonymous):

prove the slopes of the parallel sides are equal

OpenStudy (anonymous):

\[\frac{2--2}{7-3}=\frac{8-4}{5-1}=1\]

OpenStudy (anonymous):

i still dont understand how you got this

OpenStudy (anonymous):

you have to understand what your problem is saying. Its wants you to basically prove that the figure is a parrallegoram. Now it tell you that a parrallogram is only a parallegoram if if has two parrallel sides. (i assume you understand what parrallel lines are ). Now for two line to be parrallel they have to have the same slope. For that reason you are given those 4 ordered pairs. You have to determine from those orderd pairs that the slope is the same

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhh now thats what i call an explanation, nice one.

OpenStudy (anonymous):

you just got schooled

OpenStudy (anonymous):

excuse me?

OpenStudy (anonymous):

you just got schooled

OpenStudy (anonymous):

are you being legit? We're in the 21 century not in renaissance , talk ENGLISH please.

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

hey come back

OpenStudy (anonymous):

i see my explanation was rather vague, apologies

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

plot your four points, label the letters then find the slopes of the non horizontal lines they should be equal if it is a true parallelogram

OpenStudy (anonymous):

yes i figured thats how it was supposed to bee :) thanks anyway

OpenStudy (anonymous):

i am talking english